can you help me with this word problem..??
Roast Beef has 25g or protein and 11g of calcium per serving. A serving of mashed potatoes has 2g of protein and 25g of calcium. How many servings of each are needed to supply exactly 29g of protein and 61g of calcium?
Thanks so much!

Answer by ankor@dixie-net.com(22740)

Roast Beef has 25g or protein and 11g of calcium per serving.
A serving of mashed potatoes has 2g of protein and 25g of calcium.
How many servings of each are needed to supply exactly 29g of protein and 61g of calcium?
:
let b = servings of beef
let p = servings of potatoes
:
Write two equations, one for protein and one for calcium
:
Protein
25b + 2p = 29
:
Calcium
11b + 25p = 61
:
Multiply the 1st equation by 25, multiply the 2nd by 2
625b + 50p = 725
22b + 50p = 122
-------------------subtraction eliminates p
603b = 603
b = 1 serving of beef
:
Find p
25(1) + 2p = 29
2p = 29 - 25
2p = 4
p = 2 servings of potatoes
;
;
check solution in the calcium equation
11(1) + 25(2) = 61